# url]sites.google.com/view/mslee2

### v1

@@ Abstract
In this work, we address the challenges posed by the high nonlinearity of the Butler-Volmer (BV) equation in forward and inverse simulations of the pseudo-two-dimensional (P2D) model using the physics-informed neural network (PINN) framework. The BV equation presents significant challenges for PINNs, primarily due to the hyperbolic sine term, which renders the Hessian of the PINN loss function highly ill-conditioned. To address this issue, we introduce a bypassing term that improves numerical stability by substantially reducing the condition number of the Hessian matrix. Furthermore, the small magnitude of the ionic flux $j$ often leads to a common failure mode where PINNs converge to incorrect solutions. We demonstrate that incorporating a secondary conservation law for the solid-phase potential $\psi$ effectively prevents such convergence issues and ensures solution accuracy. The proposed methods prove effective for solving both forward and inverse problems involving the BV equation. Specifically, we achieve precise parameter estimation in inverse scenarios and reliable solution predictions for forward simulations.
@@ 1 Introduction
Lithium-ion batteries have become integral to modern technology, powering a wide range of devices from laptops and mobile phones to electric vehicles [1]. The 2019 Nobel Prize in Chemistry, awarded to the pioneers of lithium-ion battery technology, highlights its transformative impact on modern life. However, despite their widespread adoption, lithium-ion batteries face inherent challenges, including finite lifecycles, performance degradation, and safety risks such as thermal runaway and explosions [2, 3]. These limitations, alongside environmental and economic concerns, underscore the critical need for precise monitoring and effective management of battery systems. Ensuring reliability and safety while maximizing performance and lifespan is not only vital for current applications but also for advancing the next generation of energy storage technologies [4].
@@
Effective battery management relies heavily on the ability to accurately predict and understand changes in the battery state under varying operating and environmental conditions [5, 6]. In this context, physics-based model simulations have emerged as a powerful tool, enabling accurate battery state estimation without the need for extensive experimental data. These simulations play a pivotal role in solving both forward problems, such as predicting battery performance, and inverse problems, such as identifying key material properties and reaction kinetics from limited data [7, 8]. Moreover, the increasing emphasis on digital twin technologies highlights the central role of physics-based models in developing intelligent and efficient battery management systems. By enabling a deeper understanding of battery behavior, these models support not only safer and more sustainable energy storage solutions but also the optimization of battery design and functionality [9]. The pseudo-two-dimensional (P2D) model, also known as the Doyle-Fuller-Newman model, has been widely used in lithium-ion battery studies due to its ability to accurately capture the essential electrochemical processes, such as transport and reaction kinetics [10]. The P2D model divides the battery into three distinct regions: the negative electrode, the separator, and the positive electrode. It accounts for the transport of lithium ions through the electrolyte, the diffusion of lithium within solid particles in the electrodes, and the electrochemical reactions that occur at the interfaces between the electrolyte and solid particles. The P2D model consists of coupled partial differential equations (PDEs) to describe the concentration of lithium ions and the electric potentials across these regions, along with the Butler-Volmer equation to model the kinetics of the electrochemical reactions. Numerous studies have validated the P2D model’s accuracy, showing good agreement with experimental data and highlighting its effectiveness in predicting the dynamic behavior of lithium-ion batteries [11, 12].
@@
Numerical experiments play a crucial role in understanding and optimizing complex systems, including lithium-ion battery models. Over the years, numerous efforts have been made to develop and refine classical numerical methods for solving PDEs, such as finite element and finite difference methods. These methods have proven effective in solving the P2D model, providing accurate results under various conditions [13, 14, 15]. The availability of software packages in MATLAB [16], Python [17], and Julia [18] has further democratized the use of these methods, making them accessible to researchers and engineers across disciplines. However, the complexity of the underlying equations presents implementation challenges, and these classical methods often struggle with solving inverse problems due to high computational demands and sensitivity to noise. Recently, scientific machine learning (SciML) has emerged as a potential alternative method for solving both forward and inverse problems of PDEs. Especially, physics-informed neural networks (PINNs) incorporate governing physical laws directly into the loss function used for neural network learning, making them particularly well-suited for problems with limited data [19]. PINNs have shown promising results across various scientific domains [20, 21, 22]. These applications highlight PINNs’ ability to handle complex, multi-scale, and multi-physics problems by leveraging the underlying physics as a constraint during the optimization process. Given this success in other fields, there has been increasing interest in applying SciML techniques, particularly PINNs, to battery simulations. In the context of battery simulations, SciML techniques hold significant potential for capturing hidden physical phenomena, such as transport and reaction dynamics, while also enabling the estimation of critical parameters that are difficult to measure directly.
@@
Despite the notable progress in applying PINNs to various fields of physics and engineering, their applicability across diverse domains still remains underexplored, highlighting the need for further extensive research. In particular, vanilla PINNs have shown significant limitations in addressing highly non-linear phenomena, complex boundary conditions, and stiff problems. These challenges also arise in battery simulations [23]. Specifically, the vanilla PINN approach struggles to effectively solve the P2D model due to the highly nonlinear flux term, known as the Butler-Volmer (BV) equation, which describes electrochemical reactions in batteries, as well as the model’s multi-scale nature. To address these challenges, several studies have proposed modified approaches. Xue et al. [24] considered the single particle model (SPM), a simplified version of the P2D model, instead of directly solving the full P2D model, to reduce computational complexity and instability. Hassanaly et al. [25] proposed a two-step training strategy: first fitting the neural network to the single particle model (SPM), and then using the learned parameters as initial values for fitting the P2D model with a linearized BV equation. For the nonlinear BV equation, they suggested an additional division step to handle gradient exploding problems. Besides, Zubov et al. [26] studied a simplified P2D model using PINNs to test their software NeuralPDE.jl. Chen et al. [27] simulated redox flow batteries with PINNs, formulating a model with 6 governing equations and 24 boundary conditions.
@@
In this work, we address the specific challenges posed by the BV equation in P2D model simulations using PINNs. By designing and employing simplified paradigm models that replicate the inherent complexities of the BV equation, we systematically analyzed and identified critical issues that arise in battery simulations. Based on these insights, we developed novel strategies to overcome these challenges, enabling the direct solution of the full P2D model with the fully nonlinear BV equation in a single step, without relying on commonly used simplifications such as the SPM or linearized BV equations. The proposed strategies include the following. First, we introduce a bypassing term, which employs a neural network to approximate the stiffest component of the BV equation. This reduces the Hessian spectrum of the PINN loss, significantly improving the conditioning of the optimization problem and stabilizing the training process. Second, we augment the PINN loss function with a secondary conservation law for the solid-phase potential, ensuring that the predicted solution faithfully captures the nonlinear dynamics of the system and avoids convergence to incorrect solutions. Using these strategies, we achieve accurate simulation results. This demonstrates the efficacy of our approach in addressing the computational challenges associated with nonlinear battery models. Furthermore, we demonstrate the effectiveness of the proposed PINN framework in addressing inverse problems related to lithium-ion battery models, specifically the estimation of critical geometric parameters such as the total battery length and the length ratio of battery sections. Inverse simulations of this kind are notoriously challenging for classical numerical methods due to high computational costs and sensitivity to noise. By leveraging the strengths of the PINN framework and building on our robust solution to the forward problem, we successfully estimate these parameters directly within the nonlinear P2D model framework. Our method requires minimal modifications to incorporate observational data and avoids the need for simplifying assumptions or additional numerical approximations. These results highlight the capability of our approach to tackle complex inverse problems with improved reliability and computational efficiency.
@@ 2.1 Pseudo-Two-Dimensional Model
The P2D model, proposed in [10], captures the complex dynamics within a lithium-ion battery cell by modeling the transport and diffusion of lithium ions along the through-thickness of the cell, denoted as the $x$-direction. The model assumes that lithium ions move through the electrolyte in the $x$-direction, while lithium-ion diffusion within each spherical solid particle in the electrodes occurs symmetrically along the radial direction $r$. In this paper, we denote the lengths of the positive electrode, separator, and negative electrode as $L_{p}$, $L_{s}$, and $L_{n}$, respectively. Then, $x\in[0,L_{p}]=:\mathcal{D}_{p}$ corresponds to the positive electrode, $x\in[L_{p},L_{p}+L_{s}]=:\mathcal{D}_{s}$ represents the separator, and $x\in[L_{p}+L_{s},L_{p}+L_{s}+L_{n}]=:\mathcal{D}_{n}$ represents the negative electrode. For simplicity, we denote the total length as $L:=L_{p}+L_{s}+L_{n}$. For the radial direction within solid particles, the coordinate $r$ ranges from 00 to $\mathcal{R}_{p}$ in the positive electrode and from 00 to $\mathcal{R}_{n}$ in the negative electrode, where $\mathcal{R}_{p}$ and $\mathcal{R}_{n}$ denote the radii of the solid particles in each electrode, respectively.
The primary variables of the P2D model include: (1) the solid-phase lithium-ion concentration, $c_{s,i}$ ($i=p,n$), in the positive and negative electrodes; (2) the solid-phase potential, $\psi_{i}$ ($i=p,n$), in the electrodes; (3) the lithium-ion concentration, $c_{i}$ ($i=p,s,n$), in the electrolyte; and (4) the liquid-phase potential, $\phi_{i}$ ($i=p,s,n$), in the electrolyte. Here, $p$, $s$, and $n$ represent the positive electrode, separator, and negative electrode, respectively. Additional variables, such as the overpotential $\eta_{i}$ and ionic flux $j_{i}$, are also considered. A comprehensive summary of these state variables and other quantities used in the P2D model is provided in Table 1. Using these variables, the governing equations of the P2D model are formulated to describe the time evolution of the state variables as follows.
@@
{$-D_{s,i}^{\mathrm{eff}}\dfrac{\partial c_{s,i}}{\partial r}=0,\quad&\text{at }r=0,\\ -D_{s,i}^{\mathrm{eff}}\dfrac{\partial c_{s,i}}{\partial r}=j_{i}\ &\text{at}\ r=\mathcal{R}_{i},\end{aligned}\right.$
@@
The aforementioned state variables are coupled by the ionic flux $j_{i}$, which is given by the following Butler-Volmer equation:
@@
$c_{s}$ solid phase concentration $k$ electrochemical reaction rate constant
$c$ liquid phase concentration $\epsilon$ volume fraction
$\eta$ overpotential Brugg Bruggmann constant
$j$ ionic flux $\kappa$ ionic conductivity
@@
$t_{+}^{0}$ transference number of lithium ion
@@ 2.2 Physics-Informed Neural Networks
[…]
@@
3 Physics-Informed Neural Networks for the P2D model
@@
In this section, we apply PINNs to solve the P2D model. The P2D model includes two interfaces: one between the positive electrode and the separator, and the other between the separator and the negative electrode. The interfacial conditions enforce that for the liquid-phase concentration $c_{i}$ and the potential $\psi_{i}$, they and their flux are continuous at the interfaces of both the electrodes and the separator. The flux continuity conditions can cause discontinuities in the first derivatives, which makes it difficult for a single approximate network defined over the whole $x$-axis to satisfy the flux continuity. For this reason, we adopted the domain decomposition approach, where the computational domain is divided into three subdomains—$\mathcal{D}_{p}$, $\mathcal{D}_{s}$, and $\mathcal{D}_{n}$—with separate networks used to approximate each state variable within each subdomain. In other words, each of the main battery state variables, such as $c_{s,i}$, $c_{i}$, $\phi_{i}$, and $\psi_{i}$, is approximated by a separate neural network for each region $\mathcal{D}_{i}$ ($i=p,s,n$), denoted as $\widetilde{c}_{s,i}$, $\widetilde{c}_{i}$, $\widetilde{\phi}_{i}$, and $\widetilde{\psi}_{i}$, respectively. The interfacial conditions are enforced through penalty loss, which will be discussed in Section 3.3.
In PINN frameworks, one of the conventional ways to train the networks is to minimize the following loss, given as:
$\mathcal{L}_{\mathrm{PINN}}(\theta):=$ $\mathcal{L}_{\mathrm{PDE}}(\theta)+\lambda_{\mathrm{IC}}\mathcal{L}_{\mathrm{IC}}(\theta)+\lambda_{\mathrm{BC}}\mathcal{L}_{\mathrm{BC}}(\theta)+\lambda_{\mathrm{Inter}}\mathcal{L}_{\mathrm{Inter}}(\theta)$ (3)
@@
where $\theta$ represents the network parameter and each of loss function and $\lambda_{\bullet}$ are weighting parameters to be chosen appropriately. Each term in the loss function—$\mathcal{L}_{\mathrm{PDE}}(\theta)$, $\mathcal{L}_{\mathrm{IC}}(\theta)$, $\mathcal{L}_{\mathrm{BC}}(\theta)$, and $\mathcal{L}_{\mathrm{Inter}}(\theta)$—is conventionally defined to enforce the PDEs, boundary/initial conditions, and the interface conditions into the networks. Specifically, the PDE residual loss $\mathcal{L}_{\mathrm{PDE}}(\theta)$ can be split into four parts, corresponding to the main variables: $\mathcal{L}_{\mathrm{c_{s},PDE}}(\theta)$, $\mathcal{L}_{\mathrm{c,PDE}}(\theta)$, $\mathcal{L}_{\mathrm{\phi,PDE}}(\theta)$, and $\mathcal{L}_{\mathrm{\psi,PDE}}(\theta)$. For example, the PDE residual loss $\mathcal{L}_{\mathrm{c,PDE}}(\theta)$ for the liquid concentration $c$ is given as
[…]
where the approximate ionic flux $\widetilde{j}_{i}$ is algebraically computed by
$\widetilde{j}_{i}=2k\left(c_{s,\mathrm{max}}-\widetilde{c}_{s,i,\mathrm{surf}}\right)^{0.5}\widetilde{c}_{s,i,\mathrm{surf}}^{0.5}\widetilde{c}_{i}^{0.5}\sinh\left(\frac{F}{2RT}\left(\widetilde{\phi}_{i}-\widetilde{\psi}_{i}-U(\widetilde{c}_{s,i}/c_{\mathrm{s,max}})\right)\right),$
according to the original relation (1). The PDE residual losses for the other state variables are given similarly. Likewise, the initial, boundary, and interface condition losses can also be split into four parts. For instance, the losses for the liquid-phase concentration $c_{i}$ are given as follows:
[…]
@@
$\mathcal{L}_{\mathrm{c,BC}}(\theta)=$ $\int_{[0,\mathcal{T}]}\left|D_{p}^{\mathrm{eff}}\frac{\partial\widetilde{c}_{p}}{\partial x}(t,0)\right|^{2}+\left|D_{n}^{\mathrm{eff}}\frac{\partial\widetilde{c}_{n}}{\partial x}(t,L)\right|^{2}dt,$

### v2

# Forward and Inverse Simulation of Pseudo-Two-Dimensional Model of Lithium-Ion Batteries Using Neural Networks

@@ Abstract
In this work, we address the challenges posed by the high nonlinearity of the Butler–Volmer (BV) equation in forward and inverse simulations of the pseudo-two-dimensional (P2D) model using the physics-informed neural network (PINN) framework. The BV equation presents significant challenges for PINNs, primarily due to the hyperbolic sine term, which renders the Hessian of the PINN loss function highly ill-conditioned. To address this issue, we introduce a bypassing term that improves numerical stability by substantially reducing the condition number of the Hessian matrix. Furthermore, the small magnitude of the ionic flux $j$ often leads to a common failure mode where PINNs converge to incorrect solutions. We demonstrate that incorporating a secondary conservation law for the solid-phase potential $\psi$ effectively prevents such convergence issues and ensures solution accuracy. The proposed methods prove effective for solving both forward and inverse problems involving the BV equation. Specifically, we achieve precise parameter estimation in inverse scenarios and reliable solution predictions for forward simulations.
@@ 1 Introduction
Lithium-ion (Li-ion) batteries have become integral to modern technology, powering a wide range of devices from laptops and mobile phones to electric vehicles [1]. The 2019 Nobel Prize in Chemistry, awarded to the pioneers of Li-ion battery technology, highlights its transformative impact on modern life. However, despite their widespread adoption, Li-ion batteries face inherent challenges, including finite lifecycles, performance degradation, and safety risks such as thermal runaway and explosions [2, 3]. These limitations, alongside environmental and economic concerns, underscore the critical need for precise monitoring and effective management of battery systems. Ensuring reliability and safety while maximizing performance and lifespan is not only vital for current applications but also for advancing the next generation of energy storage technologies [4].
@@
Effective battery management relies heavily on the ability to accurately predict and understand changes in the battery state under varying operating and environmental conditions [5, 6]. In this context, physics-based model simulations have emerged as a powerful tool, enabling accurate battery state estimation without the need for extensive experimental data. These simulations play a pivotal role in solving both forward problems, such as predicting battery performance, and inverse problems, such as identifying key material properties and reaction kinetics from limited data [7, 8]. Moreover, the increasing emphasis on digital twin technologies highlights the central role of physics-based models in developing intelligent and efficient battery management systems. By enabling a deeper understanding of battery behavior, these models support not only safer and more sustainable energy storage solutions but also the optimization of battery design and functionality [9]. The pseudo-two-dimensional (P2D) model, also known as the Doyle–Fuller–Newman model, has been widely used in Li-ion battery studies due to its ability to accurately capture the essential electrochemical processes, such as transport and reaction kinetics [10]. The P2D model divides the battery into three distinct regions: the negative electrode, the separator, and the positive electrode. It accounts for the transport of lithium ions through the electrolyte, the diffusion of lithium within solid particles in the electrodes, and the electrochemical reactions that occur at the interfaces between the electrolyte and solid particles. The P2D model consists of coupled partial differential equations (PDEs) to describe the concentration of lithium ions and the electric potentials across these regions, along with the Butler–Volmer (BV) equation to model the kinetics of the electrochemical reactions. Numerous studies have validated the P2D model’s accuracy, showing good agreement with experimental data and highlighting its effectiveness in predicting the dynamic behavior of Li-ion batteries [11, 12].
@@
Numerical experiments play a crucial role in understanding and optimizing complex systems, including Li-ion battery models. Over the years, numerous efforts have been made to develop and refine classical numerical methods for solving PDEs, such as finite element and finite difference methods. These methods have proven effective in solving the P2D model, providing accurate results under various conditions [13, 14, 15]. The availability of software packages in MATLAB [16], Python [17], and Julia [18] has further democratized the use of these methods, making them accessible to researchers and engineers across disciplines. However, the complexity of the underlying equations presents implementation challenges, and these classical methods often struggle with solving inverse problems due to high computational demands and sensitivity to noise. Recently, scientific machine learning (SciML) has emerged as a potential alternative method for solving both forward and inverse problems of PDEs. Especially, physics-informed neural networks (PINNs) incorporate governing physical laws directly into the loss function used for neural network learning, making them particularly well-suited for problems with limited data [19]. PINNs have shown promising results across various scientific domains [20, 21, 22]. These applications highlight PINNs’ ability to handle complex, multi-scale, and multi-physics problems by leveraging the underlying physics as a constraint during the optimization process. Given this success in other fields, there has been increasing interest in applying SciML techniques, particularly PINNs, to battery simulations. In the context of battery simulations, SciML techniques hold significant potential for capturing hidden physical phenomena, such as transport and reaction dynamics, while also enabling the estimation of critical parameters that are difficult to measure directly.
@@
Despite the notable progress in applying PINNs to various fields of physics and engineering, their applicability across diverse domains still remains underexplored, highlighting the need for further extensive research. In particular, vanilla PINNs have shown significant limitations in addressing highly non-linear phenomena, complex boundary conditions, and stiff problems. These challenges also arise in battery simulations [23]. Specifically, the vanilla PINN approach struggles to effectively solve the P2D model due to the highly nonlinear flux term, known as the BV equation, which describes electrochemical reactions in batteries, as well as the model’s multi-scale nature. To address these challenges, several studies have proposed modified approaches. Xue et al. [24] considered the single particle model (SPM), a simplified version of the P2D model, instead of directly solving the full P2D model, to reduce computational complexity and instability. Hassanaly et al. [25] proposed a two-step training strategy: first fitting the neural network to the single particle model (SPM), and then using the learned parameters as initial values for fitting the P2D model with a linearized BV equation. For the nonlinear BV equation, they suggested an additional division step to handle gradient exploding problems. Besides, Zubov et al. [26] studied a simplified P2D model using PINNs to test their software NeuralPDE.jl. Chen et al. [27] simulated redox flow batteries with PINNs, formulating a model with 6 governing equations and 24 boundary conditions.
@@
In this work, we address the specific challenges posed by the BV equation in P2D model simulations using PINNs. By designing and employing simplified toy models that replicate the inherent complexities of the BV equation, we systematically analyzed and identified critical issues that arise in battery simulations. Based on these insights, we developed novel strategies to overcome these challenges, enabling the direct solution of the full P2D model with the fully nonlinear BV equation in a single step, without relying on commonly used simplifications such as the SPM or linearized BV equations. The proposed strategies include the following. First, we introduce a bypassing term, which employs a neural network to approximate the stiffest component of the BV equation. This reduces the Hessian spectrum of the PINN loss, significantly improving the conditioning of the optimization problem and stabilizing the training process. Second, we augment the PINN loss function with a secondary conservation law for the solid-phase potential, ensuring that the predicted solution faithfully captures the nonlinear dynamics of the system and avoids convergence to incorrect solutions. Using these strategies, we achieve accurate simulation results. This demonstrates the efficacy of our approach in addressing the computational challenges associated with nonlinear battery models. Furthermore, we demonstrate the effectiveness of the proposed PINN framework in addressing inverse problems related to Li-ion battery models, specifically the estimation of critical geometric parameters such as the total battery length and the length ratio of battery sections. Inverse simulations of this kind are notoriously challenging for classical numerical methods due to high computational costs and sensitivity to noise. By leveraging the strengths of the PINN framework and building on our robust solution to the forward problem, we successfully estimate these parameters directly within the nonlinear P2D model framework. Our method requires minimal modifications to incorporate observational data and avoids the need for simplifying assumptions or additional numerical approximations. These results highlight the capability of our approach to tackle complex inverse problems with improved reliability and computational efficiency.
@@ 2.1 Pseudo-two-dimensional model
The P2D model, proposed in [10], captures the complex dynamics within a Li-ion battery cell by modeling the transport and diffusion of lithium ions along the through-thickness of the cell, denoted as the $x$-direction. The model assumes that lithium ions move through the electrolyte in the $x$-direction, while Li-ion diffusion within each spherical solid particle in the electrodes occurs symmetrically along the radial direction $r$. In this paper, we denote the lengths of the positive electrode, separator, and negative electrode as $L_{p}$, $L_{s}$, and $L_{n}$, respectively. Then, $x\in[0,L_{p}]=:\mathcal{D}_{p}$ corresponds to the positive electrode, $x\in[L_{p},L_{p}+L_{s}]=:\mathcal{D}_{s}$ represents the separator, and $x\in[L_{p}+L_{s},L_{p}+L_{s}+L_{n}]=:\mathcal{D}_{n}$ represents the negative electrode. For simplicity, we denote the total length as $L:=L_{p}+L_{s}+L_{n}$. For the radial direction within solid particles, the coordinate $r$ ranges from 00 to $\mathcal{R}_{p}$ in the positive electrode and from 00 to $\mathcal{R}_{n}$ in the negative electrode, where $\mathcal{R}_{p}$ and $\mathcal{R}_{n}$ denote the radii of the solid particles in each electrode, respectively. For descriptive purposes, we provide a schematic representation of the P2D model in Figure 1.
Figure 1: A schematic of the P2D model.
The primary variables of the P2D model include: (1) the solid-phase Li-ion concentration, $c_{s,i}$ ($i=p,n$), in the positive and negative electrodes; (2) the solid-phase potential, $\psi_{i}$ ($i=p,n$), in the electrodes; (3) the Li-ion concentration, $c_{i}$ ($i=p,s,n$), in the electrolyte; and (4) the liquid-phase potential, $\phi_{i}$ ($i=p,s,n$), in the electrolyte. Here, $p$, $s$, and $n$ represent the positive electrode, separator, and negative electrode, respectively. Additional variables, such as the overpotential $\eta_{i}$ and ionic flux $j_{i}$, are also considered. A comprehensive summary of these state variables and other quantities used in the P2D model is provided in Table 1. Using these variables, the governing equations of the P2D model are formulated to describe the time evolution of the state variables as follows.
@@
{$-D_{s,i}^{\mathrm{eff}}\dfrac{\partial c_{s,i}}{\partial r}=0,\quad&\text{at }r=0,\\ -D_{s,i}^{\mathrm{eff}}\dfrac{\partial c_{s,i}}{\partial r}=j_{i},\ &\text{at}\ r=\mathcal{R}_{i}.\end{aligned}\right.$
@@
The aforementioned state variables are coupled by the ionic flux $j_{i}$, which is given by the following BV equation:
@@
$c_{s}$ solid phase concentration $c_{s,\mathrm{max}}$ maximum value for $c_{s}$
$c$ liquid phase concentration $c_{\mathrm{ref}}$ reference value for $c$
$j$ ionic flux $k$ electrochemical reaction rate constant
@@
$\epsilon$ volume fraction $t_{+}^{0}$ transference number of Li-ion
$\eta$ overpotential Brugg Bruggmann constant
$\kappa$ ionic conductivity $\bullet^{\mathrm{eff}}$ effective coefficient: ×$\bullet\times\epsilon^{\mathrm{Brugg}}$
@@ 2.2 Physics-informed neural networks
$\mathcal{L}_{\mathrm{PINN}}(\theta)=\int_{\Omega}\left\lvert\mathcal{N}[u_{\theta}](x)-f(x)\right\rvert_{2}^{2}\mathrm{d}x+\lambda\int_{\partial\Omega}\left\lvert\mathcal{B}[u_{\theta}](s)-g(s)\right\rvert_{2}^{2}\mathrm{d}\sigma(s).$ (2)
@@
3 Conventional application of PINNs to the P2D model
In this section, we first provide a complete presentation of how PINNs are typically applied to PDEs, which we refer to as the “vanilla PINN” approach. We then discuss some issues that can appears in applying the vanilla PINNs and address the issues with existing methods.
3.1 Vanilla PINNs for the P2D model
@@
In this subsection, we present a complete explanation of a vanilla PINN approach to solve the P2D model. The P2D model includes two interfaces: one between the positive electrode and the separator, and the other between the separator and the negative electrode. The interfacial conditions enforce that for the liquid-phase concentration $c_{i}$ and the potential $\psi_{i}$, they and their flux are continuous at the interfaces of both the electrodes and the separator. The flux continuity conditions can cause discontinuities in the first derivatives, which makes it difficult for a single approximate network defined over the whole $x$-axis to satisfy the flux continuity. To address this, we adopt a domain decomposition approach, dividing the $x$-axis into three subdomains $\mathcal{D}_{p}$, $\mathcal{D}_{s}$, and $\mathcal{D}_{n}$. We then assign a separate neural network to approximate each primary state variable (e.g. $c_{s,i},c_{i},\phi_{i},\psi_{i}$) in each subdomain. In other words, each region $i\in\{p,s,n\}$ is handled by its own network, denoted by
$\widetilde{c}_{s,i},\quad\widetilde{c}_{i},\quad\widetilde{\phi}_{i},\quad\widetilde{\psi}_{i}.$
With such approximations for the primary variables by neural networks, the ionic flux $j_{i}$ is approximated by
$\widetilde{j}_{i}=2k\left(c_{s,\mathrm{max}}-\widetilde{c}_{s,i,\mathrm{surf}}\right)^{0.5}\widetilde{c}_{s,i,\mathrm{surf}}^{0.5}\widetilde{c}_{i}^{0.5}\sinh\left(\frac{F}{2RT}\left(\widetilde{\phi}_{i}-\widetilde{\psi}_{i}-U(\widetilde{c}_{s,i}/c_{\mathrm{s,max}})\right)\right),$ (3)
according to the original relation (1). Then, the PINN loss function is given as follows:
$\mathcal{L}_{\mathrm{PINN}}(\theta):=\mathcal{L}_{\mathrm{PDE}}(\theta)+\lambda_{\mathrm{IC}}\mathcal{L}_{\mathrm{IC}}(\theta)+\lambda_{\mathrm{BC}}\mathcal{L}_{\mathrm{BC}}(\theta)+\lambda_{\mathrm{Inter}}\mathcal{L}_{\mathrm{Inter}}(\theta),$ (4)
@@
where $\theta$ represents the network parameter and each of the loss function terms and $\lambda_{\bullet}$ are weighting parameters to be chosen appropriately. Each term in the loss function—$\mathcal{L}_{\mathrm{PDE}}(\theta)$, $\mathcal{L}_{\mathrm{IC}}(\theta)$, $\mathcal{L}_{\mathrm{BC}}(\theta)$, and $\mathcal{L}_{\mathrm{Inter}}(\theta)$—is conventionally defined to enforce the PDEs, boundary/initial conditions, and the interface conditions into the networks.
Specifically, the PDE residual loss $\mathcal{L}_{\mathrm{PDE}}(\theta)$ can be split into four parts, corresponding to the main variables: $\mathcal{L}_{\mathrm{c_{s},PDE}}(\theta)$, $\mathcal{L}_{\mathrm{c,PDE}}(\theta)$, $\mathcal{L}_{\mathrm{\phi,PDE}}(\theta)$, and $\mathcal{L}_{\mathrm{\psi,PDE}}(\theta)$. For example, the PDE residual loss $\mathcal{L}_{\mathrm{c,PDE}}(\theta)$ for the liquid concentration $c$ is given as
@@
$\begin{split}\mathcal{L}_{\mathrm{c,PDE}}(\theta)=\sum_{i\in\{p,s,n\}}\int_{[0,\mathcal{T}]\times\mathcal{D}_{i}\times[0,\mathcal{R}_{i}]}\left|\epsilon_{i}\frac{\partial\widetilde{c}_{i}}{\partial t}-\frac{\partial}{\partial x}\left(D^{\mathrm{eff}}\frac{\partial\widetilde{c}_{i}}{\partial x}\right)-(1-t_{+}^{0})a_{i}\widetilde{j}_{i}\right|^{2}\ \mathrm{d}t\mathrm{d}x\mathrm{d}r,\end{split}.$
The PDE residual losses for the other state variables are given similarly. Likewise, the initial, boundary, and interface condition losses can also be split into four parts. For instance, the losses for the liquid-phase concentration $c_{i}$ are given as follows:
$\mathcal{L}_{\mathrm{c,IC}}(\theta)=$ $\sum_{i=p,s,n}\int_{\mathcal{D}_{i}}\left|\widetilde{c}_{i}(0,x)-c_{i}^{0}(x)\right|^{2}\mathrm{d}x,$
$\mathcal{L}_{\mathrm{c,BC}}(\theta)=$ $\int_{[0,\mathcal{T}]}\left|D_{p}^{\mathrm{eff}}\frac{\partial\widetilde{c}_{p}}{\partial x}(t,0)\right|^{2}+\left|D_{n}^{\mathrm{eff}}\frac{\partial\widetilde{c}_{n}}{\partial x}(t,L)\right|^{2}\mathrm{d}t,$